You are given the production function: Q(K,L)=10KαLβ

Does the above function exhibit increasing, decreasing or constant return to scale? Illustrate why and explain what this means.


1
Expert's answer
2021-05-16T17:46:25-0400

Production function is given as;

"Q(K,L)=10K^\u03b1L^\u03b2"

For the determination of the type of return to scale we have to make certain changes in inputs, let say both the inputs are changed by x.

Then,

"Q(xK,xL)=10(xK)^\u03b1(xL)^\u03b2"

"= 10 (x)^\u03b1 + \u03b2 (K^\u03b1L^\u03b2)"

"= (x)^\u03b1 + \u03b2(10K^\u03b1L^\u03b2)"

"= (x)^\u03b1 + \u03b2 Q(K,L)"

Here return to scale depend upon the α + β

We could have three cases because the value of α, β are not specified

If (α + β) = 1 constant return to scale

If (α + β) > 1 positive return to scale

If (α + β) < 1 negative return to scale